OpenStudy (mgriffin16):

Harold left town driving east at 55 mph. Carol left 2 hours later driving the same route at 65 mph. How many hours after Harold left will it take Carol to catch up to Harold?

OpenStudy (welshfella):

Hint: they will have both travelled the same distance when Carol catches up with Harold. Use the formula V = d / t (speed = distance / time)

OpenStudy (mgriffin16):

so would i do 65x120(since that is two hours) and then divide by 55?

OpenStudy (welshfella):

let t = time after harold left when carol caught up then t - 2 is the time that carol took to catch him up they both travelled a distance of d miles for Carol 65 = d / (t-2) so d = 65(t - 2) for harold: 55 = d / t so d = 55t therfore 55t = 65(t - 2) solve this for t and you have your answer

OpenStudy (welshfella):

Do you follow the reasoning?